# Joint Design of Radar Receive Filter and Unimodular ISAC Waveform with Sidelobe Level Control

### v1

@@ II-B Signal Model for Radar Imaging
As shown in Fig. 2, after removing the CP, the received echos for radar imaging at the $i$-th range bin, $\mathbf{D}_{(i)}\in\mathbb{C}^{N_{R}\times T}$, are obtained from the $t_{i+L}$-th time slot to the $t_{i+L+T-1}$-th time slot, where $t_{i+L}$ is the starting time slot of the $i$-th range bin, and $\mathbf{D}_{(i)}\in\mathbb{C}^{N_{R}\times T}$ can be expressed as [11, eq. (13)]
@@
where $\mathbf{X}$ is the transmitted ISAC waveform defined above, $\mathbf{Z}_{\text{R}}$ is the additive Gaussian noise matrix with zero mean and covariance matrix $\mathbf{R}_{\text{s}}$, $\{\mathbf{a}(\theta_{p_{(i)}})\in\mathbb{C}^{N_{R}\times 1}\}_{p_{(i)}=1}^{P_{(i)}}$ and $\{\mathbf{v}(\theta_{p_{(i)}})\in\mathbb{C}^{N_{T}\times 1}\}_{p_{(i)}=1}^{P_{(i)}}$ are the radar receive and transmit steering vectors for the $P_{(i)}$ targets in the $i$-th range bin, respectively, and $\{h_{p_{(i)}}\in\mathbb{C}\}_{p_{(i)}=1}^{P_{(i)}}$ are complex amplitudes proportional to the RCS of these $P_{(i)}$ targets. The parameters with subscript $p_{k_{(i)}}^{\prime}$ are defined for the $P_{k_{(i)}}^{\prime}$ scatterers at the $k_{(i)}$-th range bin with the same meanings, the temporal shifting matrix $\mathbf{J}_{k}$ is defined as
@@
Under the signal models (6) and (7), we consider optimizing two metrics to improve radar imaging performance. One is the radar receive beampattern MSE, defined as $\|\mathbf{X}\mathbf{F}^{\text{H}}-\mathbf{R}_{\text{d}}\|_{\text{F}}$, measuring how close the receive beampattern approximates the desired one, where $\mathbf{R}_{\text{d}}$ is the desired spatial response determined by the prior knowledge about the targets [26]. Another metric is the maximum sidelobe level $\max_{k\in\mathbf{\Omega}_{K}}\{\|\mathbf{X}\mathbf{J}_{k}\mathbf{F}^{\text{H}}\|_{\text{F}}\}$, which is required to be less than a preset level $\xi$ for all $k\in\mathbf{\Omega}_{K}$ to reduce the clutters from neighboring range bins.
@@ II-C Problem Formulation
Based on the above discussions, the joint design problem of the radar receive filter and the ISAC waveform is formulated to minimize a weighted sum of the receive radar beampattern MSE and MUI at communication receivers, under the constraints of unimodular transmit sequence and the full power limit of the radar receive filter [^2] We adopt the full-power radar receive filter to ensure that the mainlobe energy remains as high as possible while suppressing the sidelobe level. Although this can be achieved by maximizing the sidelobe-to-mainlobe ratio, it leads to a fractional programming problem, increasing the complexity of the solution. In future work, we will consider using the sidelobe-to-mainlobe ratio as an optimization objective or constraint.. The optimization problem is presented below:
@@
where $f(\mathbf{F},\mathbf{X})=\alpha\|\mathbf{X}\mathbf{F}^{\text{H}}-\mathbf{R}_{\text{d}}\|_{\text{F}}^{2}+(1-\alpha)\|\mathbf{H}\mathbf{x}-\mathbf{s}\|_{2}^{2}$.
Solving problem (8a) is not straightforward due to the presence of both convex PSL constraints and nonconvex constraints on the receive filter and transmit sequence. In the next section, we propose an inexact ALM algorithm for solving problem (8a) efficiently. “Inexact” here refers to solving the subproblems in the ALM algorithm inexactly. By doing so, we can significantly reduce the computational cost of solving the ALM subproblems without sacrificing the solution quality.
@@ III Proposed Approach
In this section, we propose an inexact ALM algorithm for solving problem (8a). More specifically, we first introduce the main idea of the proposed inexact ALM algorithm in Section III-A. Then, we propose a scheme for finding a feasible point of problem (8a) in Section III-B, which plays an important role in guaranteeing the convergence of the proposed algorithm to a feasible stationary point. We custom-design a BSUM algorithm for efficiently solving the ALM subproblem in Section III-C. Finally, we analyze the convergence of the proposed algorithm in Section III-D.
@@ III-A Framework of Proposed ALM Algorithm
The inexact ALM algorithm is proposed to solve problem (8a). Its basic idea is to decompose the original optimization problem into smaller and more manageable subproblems. At each iteration, the algorithm updates primal variables associated with each ALM subproblem with a penalty term for penalizing the violation of the constraints in the original problem, followed by an update of the dual variables. Unlike the classic ALM algorithm, which solves each subproblem exactly, we solve each subproblem inexactly at each iteration for computational efficiency.
To present the inexact ALM algorithm, we first reformulate problem (8a) by introducing auxiliary variables $\bm{C}=(\mathbf{C}_{k})_{k\in\mathbf{\Omega}_{K}}$ with $\mathbf{C}_{k}\in\mathbb{C}^{N_{T}\times N_{R}}$ for each $k\in\mathbf{\Omega}_{K}$. Then problem (8a) becomes
@@
+$\qquad\quad\qquad\qquad+\mathcal{R}\{tr[{\mathbf{U}_{k}}^{\text{H}}(\mathbf{X}\mathbf{J}_{k}\mathbf{F}^{\text{H}}-\mathbf{C}_{k})]\}\Big{]}.$ (10)
@@ III-A1 Initialization
$\mathcal{S}_{0}$ =$=\left\{(\bm{C},\mathbf{F},\mathbf{X})\mid\mathbf{X}\mathbf{J}_{k}\mathbf{F}^{\text{H}}=\mathbf{C}_{k},~{}\forall~{}k\in\mathbf{\Omega}_{k}\right\},$
$\mathcal{S}$ =$=\{(\bm{C},\mathbf{F},\mathbf{X})\mid\mathbf{C}_{k}\in\mathcal{S}_{\text{C}},~{}\forall~{}k\in\mathbf{\Omega}_{k};\mathbf{F}\in\mathcal{S}_{\text{F}};\mathbf{X}\in\mathcal{S}_{\text{X}}\},$
@@
$\mathcal{S}_{\text{C}}$ =$=\left\{\mathbf{C}\in\mathbb{C}^{N_{T}\times N_{T}}\mid\|\mathbf{C}\|_{\text{F}}\leq\xi\right\},$ (11)
$\mathcal{S}_{\text{F}}$ =$=\{\mathbf{F}\in\mathbb{C}^{N_{T}\times T}\mid\|\mathbf{F}\|_{\text{F}}^{2}=P_{\text{F}}\},$
$\mathcal{S}_{\text{X}}$ =$=\{\mathbf{X}\in\mathbb{C}^{N_{T}\times T}\mid|x_{ij}|=P_{x},\forall~{}x_{ij}\in\mathbf{X}\}.$
@@
The upperboundness of the objective function and the ALF are important in proving the convergence of the proposed ALM algorithm to a feasible stationary point. In our proposed ALM algorithm, the penalty parameter is adaptively updated. The adaptive update of the penalty parameter enables a fast convergence but also makes the convergence analysis difficult. The upperboundness here can overcome the difficulty in the convergence analysis and guarantees that any limit point of the sequence generated by the proposed algorithm is always feasible. More details on the convergence analysis will be presented in Section III-D.
@@ III-A2 Solving the ALM Subproblem Inexactly
The notations $\partial\mathbb{I}_{\mathcal{S}_{\text{F}}}(\mathbf{F})$, $\partial\mathbb{I}_{\mathcal{S}_{\text{X}}}(\mathbf{X})$, and $\partial\mathbb{I}_{\mathcal{S}_{\text{C}}}(\mathbf{C}_{k})$ are subdifferentials of indicator functions; see [29] and [30, Corollary 8.20] for more details. The $\varepsilon$-stationary point defined in Definition 1 reduces to the standard stationary point when setting $\varepsilon=0$.
@@ III-A3 Updating Lagrange Multipliers
In the above, $\mathbf{U}_{k}$ is first updated through the standard updating rule [28] in (15a); then, each element in $\mathbf{U}_{k}$ is projected onto an interval $[-u_{\max},u_{\max}]$ with $u_{\max}>0$ being a preset constant. The projection onto the bounded set in (15a) guarantees that the Lagrange multiplies are uniformly bounded and plays a central role in guaranteeing the convergence of the proposed algorithm.
@@ III-A4 Updating the Penalty Parameter
To accelerate the convergence speed, we propose an adaptive update rule for the penalty parameter. Denote the violation of the constraints after the $\ell$-th iteration as $v^{(\ell+1)}$, i.e.,
@@
where $\delta>0$ and $\gamma>1$. The update rule in (17) will increase the penalty parameter if the violation at the current iteration is not reduced sufficiently compared with that in the previous one.
Algorithm 1 An Inexact ALM Algorithm for Problem (8a)
@@
2:     Obtain an $\varepsilon^{(\ell)}$-stationary point $(\bm{z}^{(\ell+1)},\bm{U}^{(\ell+1)})$ by Algorithm 2 (details provided in Section III-C);
@@
6:  $(\bm{z}^{(\ell+1)},\bm{U}^{(\ell+1)})$.
The proposed inexact ALM algorithm for solving problem (LABEL:prob:FX_with_C) is summarized in Algorithm 1. By setting a positive sequence $\{\varepsilon^{(\ell)}\}_{\ell\geq 0}$ with $\varepsilon^{(\ell)}\rightarrow 0$ as $\ell\rightarrow+\infty$, the limit point of the sequence generated by Algorithm 1, $(\bm{z}^{(\infty)},\bm{U}^{(\infty)})$, is a stationary point of the ALF in (III-A). Besides, since $\mathbf{X}^{(\infty)}\mathbf{J}_{k}{{\mathbf{F}}^{(\infty)}}^{\text{H}}=\mathbf{C}_{k}^{(\infty)}$ holds for all $k\in\mathbf{\Omega}_{K}$, which will be proved further ahead, then $\bm{z}^{(\infty)}$ is a feasible stationary point of problem (LABEL:prob:FX_with_C). According to [31], any feasible stationary point of (LABEL:prob:FX_with_C) is also a feasible stationary point of (8a), i.e., $(\mathbf{F}^{(\infty)},\mathbf{X}^{(\infty)})$ is a feasible stationary point of (8a). The detailed convergence analysis will be given in Section III-D.
@@ III-B2 Minimizing the Sidelobe Level Regarding $\mathbf{F}$
Suppose that the obtained solution of problem (19) is not feasible, we further minimize the sidelobe level with respect to $\mathbf{F}$ in this part. By denoting $\mathbf{\Phi}=\sum_{k\in\mathbf{\Omega}_{K}}\mathbf{J}_{k}^{\text{H}}\mathbf{X}^{\text{H}}\mathbf{X}\mathbf{J}_{k}$ and its singular value decomposition (SVD) as $\mathbf{\Phi}=\mathbf{U}_{\Phi}\mathbf{\Sigma}_{\Phi}\mathbf{U}_{\Phi}^{\text{H}}$ (assuming that the singular values are arranged in descending order), the ISL becomes
@@ III-C BSUM Algorithm for Inexactly Solving Subproblem (13)
Denote $\bm{z}^{(t)}=(\bm{C}^{(t)},\mathbf{F}^{(t)},\mathbf{X}^{(t)})$. Next we present the update of each block of variables in the BSUM method.
@@ III-C1 Solving for Auxiliary Variables
where $\tilde{\mathbf{C}}_{k}=\frac{1}{\rho^{(\ell)}+\beta}\left(\rho^{(\ell)}\mathbf{X}^{(t)}\mathbf{J}_{k}{\mathbf{F}^{(t)}}^{\text{H}}+\mathbf{U}_{k}^{(\ell)}+\beta\mathbf{C}_{k}\right)$.

### v3

@@ II-B Signal Model for Radar Imaging
As shown in Fig. 2, after removing the CP, the received echos for radar imaging at the $i$-th range bin, $\mathbf{D}_{(i)}\in\mathbb{C}^{N_{R}\times T}$, are obtained from the $t_{i+L}$-th time slot to the $t_{i+L+T-1}$-th time slot, where $t_{i+L}$ is the starting time slot of the $i$-th range bin, and $\mathbf{D}_{(i)}\in\mathbb{C}^{N_{R}\times T}$ can be expressed as [29, eq. (13)]
@@
where $\mathbf{X}$ is the transmitted ISAC waveform defined above, $\mathbf{Z}_{\text{R}}$ is the additive Gaussian noise matrix with zero mean and covariance matrix $\mathbf{R}_{\text{s}}$, $\{\mathbf{a}(\theta_{p_{(i)}})\in\mathbb{C}^{N_{R}\times 1}\}_{p_{(i)}=1}^{P_{(i)}}$ with its $n$-th element being $e^{-j2\pi n\frac{d}{\lambda}\sin(\theta_{p,i})}$, $n=0,1,\dots,N_{R}-1$, and $\{\mathbf{v}(\theta_{p_{(i)}})=[1,e^{-j2\pi\frac{d}{\lambda}\sin(\theta_{p,i})},\dots,e^{-j2\pi(N_{T}-1)\frac{d}{\lambda}\sin(\theta_{p,i})}]\in\mathbb{C}^{N_{T}\times 1}\}_{p_{(i)}=1}^{P_{(i)}}$ are the radar receive and transmit steering vectors for the $P_{(i)}$ targets in the $i$-th range bin, respectively, $d=\lambda/2$, and $\{h_{p_{(i)}}\in\mathbb{C}\}_{p_{(i)}=1}^{P_{(i)}}$ are complex amplitudes proportional to the RCS of these $P_{(i)}$ targets. The parameters with subscript $p_{k_{(i)}}^{\prime}$ are defined for the $P_{k_{(i)}}^{\prime}$ scatterers at the $k_{(i)}$-th range bin with the same meanings, the temporal shifting matrix $\mathbf{J}_{k}$ is defined as
@@
Under the signal models (6) and (7), we consider optimizing two metrics to improve radar imaging performance. One is the radar receive beampattern MSE, defined as $\|\mathbf{X}\mathbf{F}^{\text{H}}-\mathbf{R}_{\text{d}}\|_{\text{F}}$, measuring how close the receive beampattern approximates the desired one, where $\mathbf{R}_{\text{d}}$ is the desired spatial response determined by the prior knowledge about the targets [31] so that we can have stronger angle responses of the targets.For example, if we roughly know there are three targets at $10^{\circ}$, $0^{\circ}$, and $10^{\circ}$, we can solve the correlation matrix $\mathbf{R}_{\text{d}}$ through the algorithm in [34]. Otherwise, we set $\mathbf{R}_{\text{d}}$ as an identity matrix, which corresponds to an omnidirectional beampattern, if we have no knowledge about the targets. Another metric is the maximum sidelobe level $\max_{k\in\mathbf{\Omega}_{K}}\{\|\mathbf{X}\mathbf{J}_{k}\mathbf{F}^{\text{H}}\|_{\text{F}}\}$, which is required to be less than a preset level $\xi$ for all $k\in\mathbf{\Omega}_{K}$ to reduce the clutters from neighboring range bins and to ensure a clear response on the range bin of interest.
@@ II-C Problem Formulation
Based on the above discussions, the joint design problem of the radar receive filter and the ISAC waveform is formulated to minimize a weighted sum of the receive radar beampattern MSE and MUI at communication receivers under the constraints of unimodular transmit sequence and the full power limit of the radar receive filter [^2] We adopt the full-power radar receive filter to ensure that the mainlobe energy remains as high as possible while suppressing the sidelobe level. Although this can be achieved by maximizing the sidelobe-to-mainlobe ratio, it leads to a fractional programming problem, increasing the complexity of the solution. In future work, we will consider using the sidelobe-to-mainlobe ratio as an optimization objective or constraint.. The optimization problem is presented below:
@@
where $f(\mathbf{F},\mathbf{X})=\alpha\|\mathbf{X}\mathbf{F}^{\text{H}}-\mathbf{R}_{\text{d}}\|_{\text{F}}^{2}+(1-\alpha)\|\mathbf{H}\mathbf{x}-\mathbf{s}\|_{2}^{2}$, and (8d) refers to the unimodular constraint on the transmit sequence.
Solving problem (8) is not straightforward due to the presence of both convex PSL constraints and nonconvex constraints on the receive filter and transmit sequence. In the next section, we propose an inexact ALM algorithm for solving problem (8) efficiently. “Inexact” here refers to solving the subproblems in the ALM algorithm inexactly. By doing so, we can significantly reduce the computational cost of solving the ALM subproblems without sacrificing the solution quality.
@@ III Proposed Approach
In this section, we propose an inexact ALM algorithm for solving problem (8). More specifically, we first introduce the main idea of the proposed inexact ALM algorithm in Section III-A. Then, we propose a scheme for finding a feasible point of problem (8) in Section III-B, which plays an important role in guaranteeing the convergence of the proposed algorithm to a feasible stationary point. We custom-design a BSUM algorithm for efficiently solving the ALM subproblem in Section III-C. Finally, we analyze the convergence of the proposed algorithm in Section III-D.
@@ III-A Framework of Proposed ALM Algorithm
The inexact ALM algorithm is proposed to solve problem (8). Its basic idea is to decompose the original optimization problem into smaller and more manageable subproblems. At each iteration, the algorithm updates primal variables associated with each ALM subproblem with a penalty term for penalizing the violation of the constraints in the original problem, followed by an update of the dual variables. Unlike the classic ALM algorithm, which solves each subproblem exactly, we solve each subproblem inexactly at each iteration for computational efficiency.
To present the inexact ALM algorithm, we first reformulate problem (8) by introducing auxiliary variables $\bm{C}=(\mathbf{C}_{k})_{k\in\mathbf{\Omega}_{K}}$ with $\mathbf{C}_{k}\in\mathbb{C}^{N_{T}\times N_{R}}$ for each $k\in\mathbf{\Omega}_{K}$. The problem (8) becomes
@@
$\qquad\quad\qquad\qquad+\mathcal{R}\{tr[{\mathbf{U}_{k}}^{\text{H}}(\mathbf{X}\mathbf{J}_{k}\mathbf{F}^{\text{H}}-\mathbf{C}_{k})]\}\Big{]}.$ (10)
@@ III-A1 Initialization
$\mathcal{S}_{0}$ $=\left\{(\bm{C},\mathbf{F},\mathbf{X})\mid\mathbf{X}\mathbf{J}_{k}\mathbf{F}^{\text{H}}=\mathbf{C}_{k},~\forall~k\in\mathbf{\Omega}_{k}\right\},$
$\mathcal{S}$ $=\{(\bm{C},\mathbf{F},\mathbf{X})\mid\mathbf{C}_{k}\in\mathcal{S}_{\text{C}},~\forall~k\in\mathbf{\Omega}_{k};\mathbf{F}\in\mathcal{S}_{\text{F}};\mathbf{X}\in\mathcal{S}_{\text{X}}\},$
@@
$\mathcal{S}_{\text{C}}$ $=\left\{\mathbf{C}\in\mathbb{C}^{N_{T}\times N_{T}}\mid\|\mathbf{C}\|_{\text{F}}\leq\xi\right\},$ (11)
$\mathcal{S}_{\text{F}}$ $=\{\mathbf{F}\in\mathbb{C}^{N_{T}\times T}\mid\|\mathbf{F}\|_{\text{F}}^{2}=P_{\text{F}}\},$
$\mathcal{S}_{\text{X}}$ $=\{\mathbf{X}\in\mathbb{C}^{N_{T}\times T}\mid|x_{ij}|=P_{x},\forall~x_{ij}\in\mathbf{X}\}.$
@@
The upperboundness of the objective function and the ALF are important in proving the convergence of the proposed ALM algorithm to a feasible stationary point. In our proposed ALM algorithm, the penalty parameter is adaptively updated. The adaptive update of the penalty parameter enables a fast convergence but also makes the convergence analysis difficult. The upperboundness here can guarantee that any limit point of the sequence generated by the proposed algorithm is always a feasible stationary point, even when the penalty parameter becomes positive infinity. More details on the convergence analysis will be presented in Section III-D.
@@ III-A2 Solving the ALM Subproblem Inexactly
The notations $\partial\mathbb{I}_{\mathcal{S}_{\text{F}}}(\mathbf{F})$, $\partial\mathbb{I}_{\mathcal{S}_{\text{X}}}(\mathbf{X})$, and $\partial\mathbb{I}_{\mathcal{S}_{\text{C}}}(\mathbf{C}_{k})$ are subdifferentials of indicator functions; see [36] and [37, Corollary 8.20] for more details. The $\varepsilon$-stationary point defined in Definition 1 reduces to the standard stationary point when setting $\varepsilon=0$.
@@ III-A3 Updating Lagrange Multipliers
In the above, $\mathbf{U}_{k}$ is first updated through the standard updating rule [35] in (15a); then, each element in $\mathbf{U}_{k}$ is projected onto an interval $u_{\max},u_{\max}]$ with $u_{\max}>0$ being a preset constant. The projection onto the bounded set in (15a) guarantees that the Lagrange multipliers are uniformly bounded and play a central role in guaranteeing the convergence of the proposed algorithm. The boundness of the multiplier sequence in the nonconvex setting remains an open research question [36]. Readers can refer to [36] for detailed discussions.
@@ III-A4 Updating the Penalty Parameter
We propose an adaptive update rule for the penalty parameter to accelerate the convergence speed. Denote the violation of the constraints after the $\ell$-th iteration as $v^{(\ell+1)}$, i.e.,
@@
where $\delta>0$ and $\gamma>1$. The update rule in (17) will increase the penalty parameter if the violation at the current iteration is not reduced sufficiently compared with the previous one. By dynamically adjusting the penalty parameter based on the current state of constraint violations and dual variable updates, this approach helps maintain an effective balance between enforcing constraint feasibility and minimizing the objective function.
Algorithm 1 An Inexact ALM Algorithm for Problem (8)
@@
2:  Obtain an $\varepsilon^{(\ell)}$-stationary point $\bm{z}^{(\ell+1)}$ by Algorithm 2 (details provided in Section III-C);
@@
[…]
The proposed inexact ALM algorithm for solving problem (LABEL:prob:FX_with_C) is summarized in Algorithm 1. By setting a positive sequence $\{\varepsilon^{(\ell)}\}_{\ell\geq 0}$ with $\varepsilon^{(\ell)}\rightarrow 0$ as $\ell\rightarrow+\infty$, the limit point of the sequence generated by Algorithm 1, $(\bm{z}^{(\infty)},\bm{U}^{(\infty)})$, is a stationary point of the ALF in (III-A). Besides, since $\mathbf{X}^{(\infty)}\mathbf{J}_{k}{{\mathbf{F}}^{(\infty)}}^{\text{H}}=\mathbf{C}_{k}^{(\infty)}$ holds for all $k\in\mathbf{\Omega}_{K}$, which will be proved further ahead, then $\bm{z}^{(\infty)}$ is a feasible stationary point of problem (LABEL:prob:FX_with_C). According to [38], any feasible stationary point of (LABEL:prob:FX_with_C) is also a feasible stationary point of (8), i.e., $(\mathbf{F}^{(\infty)},\mathbf{X}^{(\infty)})$ is a feasible stationary point of (8). The detailed convergence analysis will be given in Section III-D.
@@ III-B2 Minimizing the Sidelobe Level Regarding $\mathbf{F}$
Suppose that the obtained solution of problem (19) is not feasible. We further minimize the sidelobe level with respect to $\mathbf{F}$ in this part. By denoting $\mathbf{\Phi}=\sum_{k\in\mathbf{\Omega}_{K}}\mathbf{J}_{k}^{\text{H}}\mathbf{X}^{\text{H}}\mathbf{X}\mathbf{J}_{k}$ and its singular value decomposition (SVD) as $\mathbf{\Phi}=\mathbf{U}_{\Phi}\mathbf{\Sigma}_{\Phi}\mathbf{U}_{\Phi}^{\text{H}}$ (assuming that the singular values are arranged in descending order), the ISL becomes
@@ III-C BSUM Algorithm for Inexactly Solving Subproblem (13)
Denote $\bm{z}^{(t)}=(\bm{C}^{(t)},\mathbf{F}^{(t)},\mathbf{X}^{(t)})$. Next, we present the update of each block of variables using the BSUM method.
@@ III-C1 Solving for Auxiliary Variables
where $\tilde{\mathbf{C}}_{k}=\frac{1}{\rho^{(\ell)}+\beta}\left(\rho^{(\ell)}\mathbf{X}^{(t)}\mathbf{J}_{k}{\mathbf{F}^{(t)}}^{\text{H}}+\mathbf{U}_{k}^{(\ell)}+\beta\mathbf{C}_{k}^{(t)}\right)$.